# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

### v2

@@ Introduction
It is not difficult to see that the Pareto set is $S^{*}=\{x\in\{0,1\}^{n}\mid|x_{B_{i}}|_{1}\in[k..n-k]\cup\{0,n\},i\in[1..m/2]\}$. Hence, the Pareto front is
@@
We recall at this point that we defined the runtime of an MOEA as the time to cover the full Pareto front $F^{*}$, that is, the first time to reach a population $P$ with $F^{*}\subseteq f(P)$.
Note that for $k>1$, not all points are Pareto optimal and not all sets of mutually non-dominated points are the subsets of the Pareto front. However, we have the following result showing that the maximum number of mutually non-dominated points is at most the size of the Pareto front. Due to the limited space, all mathematical proofs could only be sketched or had to be omitted completely. They can be found in the preprint (Zheng and Doerr 2023c).
@@
Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions such that any two are not weakly dominated (w.r.t. $m$OJZJ). Then $|P|\leq M$.
For ease of reading of the proof sketches in the following sections, we call a Pareto optimum $x$ an inner Pareto optimum if for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, we have $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$. We call $u=(u_{1},n^{\prime}+2k-u_{1},\dots,u_{m/2},n^{\prime}+2k-u_{m/2})$ and $v=(v_{1},n^{\prime}+2k-v_{1},\dots,v_{m/2},n^{\prime}+2k-v_{m/2})$ neighbors if there exists $i\in[1..m/2]$ such that $u_{i}-v_{i}=1$ and $u_{j}=v_{j}$ for all $j\neq i$.
The SMS-EMOA Can Optimize $m$OJZJ
As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential runtime to cover the full Pareto front of $m$OneMinMax, $m\geq 3$, (Zheng and Doerr 2023b), which is a special case of $m$OJZJ. In this section, we analyze the runtime of the SMS-EMOA on this problem and show that it does not encounter such problems.
Algorithm Description
The SMS-EMOA is a steady-state variant (that is, the offspring size is smaller than the parent size) of the NSGA-II. Like the NSGA-II, the SMS-EMOA works with a population of fixed size $\mu$. However, in each iteration, only one offspring $x^{\prime}$ is generated from the parent population $P_{t}$. From the $\mu+1$ individuals in the combined parent and offspring population $R_{t}=P_{t}\cup\{x^{\prime}\}$, a single individual is removed. To this aim, the SMS-EMOA like the NSGA-II uses non-dominated sorting, that is, it partitions $R_{t}$ into fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$. Different from the original NSGA-II that now uses the crowding distance as the secondary criterion for the removal of individuals, the SMS-EMOA removes the individual with the smallest hypervolume contribution in the critical front (which here always is the last front $F_{i^{*}}$).
@@
$\operatorname{HV}_{r}(S)=\mathcal{L}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}\right),$
@@
Algorithm 1 gives the pseudocode of the SMS-EMOA.
@@
5:   Use fast-non-dominated-sort() (Deb et al. 2002) to partition $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
@@
Runtime of SMS-EMOA
We now analyze the runtime of the SMS-EMOA, that is, the time until its population covers the full Pareto front of $m$OJZJ. We start by proving the following result ensuring that certain individuals survive in the population or are replaced by at least as good individuals. We formulate this result in manner more general than what we need since we expect this general version to be useful in other analyses of the SMS-EMOA.
The key to the proof is that (i) if there is a point with a rank larger than one, then all solutions with rank one (that is, in $F_{1}$) will survive, and (ii) if $F_{1}$ contains two or more solutions with same function value, then only such a solution can be removed from $F_{1}$. If the population size is large enough, by the pigeon-hole principle, one of these two cases comes true.
@@
Consider any $m$-objective optimization problem. Let $M\in\mathbb{N}$ be such that any set $S$ of pairwise non-dominated solutions having different objective values satisfies $|S|\leq M$. Consider solving this problem via the SMS-EMOA with population size $\mu\geq M$ and using a reference point $r$ such that $\operatorname{HV}_{r}(\{x\})>0$ for any individual $x\in\{0,1\}^{n}$.
@@
Recall that an inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, as defined before. While it is very likely that at least one initial solution is an inner Pareto optimum, this does not happen with probability one, and hence in the following lemma we estimate the time until the population contains an inner Pareto optimum. Since this time usually is much smaller than the time to generate all remaining Pareto optima, we do not care that this is estimate could easily be improved.
The key to the proof is to note that changing a block $x_{B_{i}}$ to a bit-string with between $k$ and $n^{\prime}-k$ ones is relatively easy. By Lemma 4, an individual with this $f_{i}$ value will remain in the population. Hence a total of at most $m$ such block changes (applied to the right individual) suffice to obtain an inner Pareto optimum.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $e\mu(mk)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population contains at least one inner Pareto optimum. Then after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
The last stage is to cover the remaining Pareto front points. The following lemma shows the runtime. The key of the proof is that we can divide the Pareto front into several levels and any individual in the $i$-th level can be generated from a point in the $(i-1)$-th level by flipping the right $k$ bits.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $eM\mu n^{k}$ iterations in expectation, the full Pareto front is covered.
Summing up the runtime of all stages from Lemmas 5 to 7, we have the following theorem for the runtime of the full coverage of the Pareto front.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $e\mu(mk)^{k}(1+\ln m)+e\mu Mn+e\mu Mn^{k}$ iterations ($\mu+e\mu(mk)^{k}(1+\ln m)+e\mu Mn+e\mu Mn^{k}$ function evaluations) in expectation, the full Pareto front is covered.
Since $k\leq n^{\prime}/2=n/m$ and $M\geq 2^{m/2}$, we easily see that both runtime expressions are $O(\mu Mn^{k})$, even when allowing $k$, $m$ and $\mu$ to depend on $n$.
Runtime of the GSEMO on $m$OJZJ
Since our arguments above can easily be extended to analyze the runtime of the GSEMO on the $m$OJZJ, we quickly do so for reasons of completeness. The GSEMO is a multiobjective counterpart of the ${(1+1)}$ EA. The initial population consists of a single randomly generated solution. In each iteration, a solution is uniformly at random picked from the population to generate an offspring via standard bit-wise mutation. If this offspring is not dominated by any solution in the population, it is added to the population and all solutions weakly dominated by it are removed. It is not difficult to see that all solutions in the population of the GSEMO are mutually non-dominated. Hence, the population size of the GSEMO is at most $M$ by Lemma 3.

### v3

@@ 3.2 Characteristics
It is not difficult to see that the Pareto set is
$S^{*}=\{x\in\{0,1\}^{n}\mid\forall i\in[1..m/2],|x_{B_{i}}|_{1}\in[k..n-k]\cup\{0,n\}\}.$
Hence, the Pareto front is
@@
What is harder to determine is the size of a largest set of incomparable, that is, pairwise not weakly dominating, individuals. One could be tempted to believe that this number is equal to the size of the Pareto front, but this is not true. [^1] In fact, in the conference version [ZD24] of this work we made this wrong claim. We sincerely apologize to the research community for this mistake. Since this number is important in the analysis of MOEAs, for example, it is the best upper bound on the population size of the GSEMO algorithm, we now discuss this question in more detail.
To this aim, let us define $\overline{M}$ as the maximum size of a set of incomparable individuals (with respect to the $m$OJZJ problem for some parameters $n$, $k$, and $m$). By definition, $\overline{M}\geq M$, since the Pareto front is witnessed by a set of incomparable solutions.
The following example shows that $\overline{M}$ can be considerably larger than $M$. Let $n$ be large, $m=4$, and $k=n^{\prime}/2$. Then $F^{*}=\{(a_{1},n^{\prime}+2k-a_{1},a_{2},n^{\prime}+2k-a_{2})\mid a_{1},a_{2}\in\{n^{\prime}/2,n^{\prime},3n^{\prime}/2\}\}$ and hence $|F^{*}|=9$. We now construct a set of pairwise incomparable solutions with much larger cardinality. For $i=1,\dots,k-1$ consider the individual $x^{(i)}=1^{i}0^{n^{\prime}-i}1^{n^{\prime}/2+i}0^{n^{\prime}/2-i}$, where we use the word notation known from theoretical computer science to describe bit-strings. In this notation, for example, $1^{i}$ is a string consisting of $i$ ones, and the concatenation of two strings $u$ and $v$ is denoted by $uv$. We have $f(x^{(i)})=(k+i,i,n^{\prime}/2-i,k+n^{\prime}/2-i)$. Consequently, for $i,j\in[1..k-1]$ with $i\neq j$, the two individuals $x^{(i)}$ and $x^{(j)}$ are incomparable. This shows that $\overline{M}\geq k-1=n/4-1$, which can be considerably larger than $M=9$.
Clearly, this example is on the extreme side. Also, noting that all our runtimes will be at least of order $n^{k}$, such large values of $k$ lead to instance that are too difficult to be solved anyway. For the more typical case that $k$ is small, the following trivial upper bound is relatively tight, as wittnessed by its closeness to the lower bound $M$.
@@
Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions pairwise incomparable solutions (w.r.t. $m$OJZJ). Then $|P|\leq\overline{M}\leq(n^{\prime}+1)^{m/2}$.
Proof.
Let $P^{\prime}$ be a set of pairwise incomparable solutions of maximal cardinality. By definition, $P^{\prime}$ contains no two solutions with identical objective value. Consequently, $|P|\leq\overline{M}=|P^{\prime}|\leq f(\{0,1\}^{n})=(n^{\prime}+1)^{m/2}$. ∎
For the ease of reading of the proofs in the following sections, we call a Pareto optimum $x$ an inner Pareto optimum if for all blocks $B_{i},i\in[1..m/2]$, we have $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$. This is equivalent to saying that $|x_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ for all $i\in[1..m/2]$. We call two objective vectors $u=(u_{1},n^{\prime}+2k-u_{1},\dots,u_{m/2},n^{\prime}+2k-u_{m/2})$ and $v=(v_{1},n^{\prime}+2k-v_{1},\dots,v_{m/2},n^{\prime}+2k-v_{m/2})$ neighbors if there exists an $i\in[1..m/2]$ such that $|u_{i}-v_{i}|=1$ and $u_{j}=v_{j}$ for all $j\neq i$.
4 The SMS-EMOA Can Optimize $m$OJZJ
As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential time to cover the full Pareto front of $m$OneMinMax, $m\geq 3$, [ZD23b], which is a special case of $m$OJZJ. In this section, we analyze the runtime of the SMS-EMOA on this problem, and more generally, the $m$OJZJ problem for $k\in[1..n^{\prime}/2]$, and show that it does not encounter such problems.
4.1 Algorithm Description
The SMS-EMOA is a steady-state variant (that is, the offspring population size is much smaller than the parent population size) of the NSGA-II. Like the NSGA-II, the SMS-EMOA works with a (parent) population of fixed size $\mu$. However, in each iteration, only one offspring $x^{\prime}$ is generated from the population $P_{t}$. From the $\mu+1$ individuals in the combined parent and offspring population $R_{t}=P_{t}\cup\{x^{\prime}\}$, a single individual is removed. To this aim, the SMS-EMOA like the NSGA-II uses non-dominated sorting, that is, it partitions $R_{t}$ into fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$. Different from the original NSGA-II, which uses the crowding distance as the secondary criterion for the removal of individuals, the SMS-EMOA removes the individual with the smallest hypervolume contribution in the critical front (which here always is the last front $F_{i^{*}}$).
@@ 4.1 Algorithm Description
$\operatorname{HV}_{r}(S)=\mathcal{L}\mathopen{}\mathclose{{}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}}\right),$
@@
Algorithm 1 gives the pseudocode of the SMS-EMOA regarded in this work. We generate the offspring by mutating a single, randomly chosen parent via bit-wise mutation with mutation rate $\frac{1}{n}$. We note that all our proofs are robust to adding crossover with constant rate $p<1$, that is, if we could generate the offspring via some crossover operator with probability $p$ and via mutation otherwise, all our runtime guarantees would remain valid except for an additional $\frac{1}{1-p}$ factor.
@@
5:   Use fast-non-dominated-sort() [DPAM02] to partition $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
@@
4.2 Runtime of SMS-EMOA
We now analyze the runtime of the SMS-EMOA, that is, the time until its population covers the full Pareto front of $m$OJZJ. We start by proving the following general result ensuring that if the population size is at least the maximum size of a set of incomparable solutions, then individuals survive in the population or are replaced by at least as good individuals. We formulate this result in manner more general than what we need since we expect this version to be useful in future analyses of the SMS-EMOA (and this happened already in [WD24]).
The key to the proof is first noting that by a simple domination argument we can concentrate on individuals in $F_{1}$. For these, we argue that (i) if the combined parent and offspring population contains an individual not in $F_{1}$, then all solutions in $F_{1}$ will survive, and (ii) if $F_{1}$ contains two or more solutions with same function value, then only such a solution can be removed from $F_{1}$. If the population size is large enough, by the pigeon-hole principle, one of these two cases is always satisfied.
@@ 4.2 Runtime of SMS-EMOA
Consider any $m$-objective optimization problem. Let $\overline{M}\in\mathbb{N}$ be such that any set $S$ of incomparable solutions satisfies $|S|\leq\overline{M}$. Consider solving this problem via the SMS-EMOA with population size $\mu\geq\overline{M}$ and using a reference point $r$ such that $\operatorname{HV}_{r}(\{x\})>0$ for any individual $x\in\{0,1\}^{n}$.
@@
Proof.
Let $t\in\mathbb{N}$ and $x\in R_{t}$. By induction, it suffices to show the lemma for $s=t+1$, that is, we now show that $P_{t+1}$ contains an individual $y$ such that $y\succeq x$. Without loss of generality, we may assume that $x$ is contained in the first front $F_{1}$ of the non-dominated sorting of $R_{t}$ – if not, then there is some $\tilde{x}\in F_{1}$ with $\tilde{x}\succ x$ and showing the existence of a $y\in P_{t+1}$ with $y\succeq\tilde{x}$ immediately gives the desired $y$ with $y\succeq x$.
Hence let us assume that $x\in F_{1}$. If $|F_{1}|\leq\mu$, then all individuals of $F_{1}$ are taken into the next generation. Consequently $y:=x\in P_{t+1}$ is the desired individual. If $|F_{1}|=\mu+1$, then by our assumption $\mu\geq\overline{M}$ and the fact that $F_{1}$ is a set of incomparable solutions, we see that $F_{1}$ contains at least two individuals with identical objective value (pigeon-hole principle). Since such individuals have a hypervolume contribution of zero, and since only such individuals have a hypervolume contribution of zero, we know that the single individual $z^{\prime}$ removed from $R_{t}$ is an individual such that there is at least one more individual $z^{\prime\prime}\in R_{t}$ with $f(z^{\prime})=f(z^{\prime\prime})$. If $z^{\prime}=x$, then $y:=z^{\prime\prime}\in P_{t+1}$ is the desired individual. If $z^{\prime}\neq x$, then $x$ survives into $P_{t+1}$ and thus trivially $y:=x$ is the desired individual. ∎
Recall that an inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i},i\in[1..m/2]$, as defined before. While it is very likely that at least one initial solution is an inner Pareto optimum, this does not happen with probability one, and hence in Lemma 6 we estimate the time until the population contains an inner Pareto optimum. Since this time usually is much smaller than the time to generate all remaining Pareto optima, we do not care that this is estimate could easily be improved.
The key to the proof is to note that changing a block $x_{B_{i}}$ to a bit-string with between $k$ and $n^{\prime}-k$ ones is relatively easy. By Lemma 4, an individual with this $f_{i}$ value will remain in the population. Hence a total of at most $m/2$ such block changes (applied to the right individual) suffice to obtain an inner Pareto optimum. In the proof, we need the following elementary observation, which extends an analogous result from [BZLQ23] for the special case that $n^{\prime}=n$.
@@
Let $k,n^{\prime},n\in\mathbb{N}$ such that $k\leq n^{\prime}/2$ and $n^{\prime}\leq n$. Then $\binom{n^{\prime}-i}{k-i}/n^{k-i}$ increases in $i\in[0..k-1]$.
Proof.
Recall that in the proof of [BZLQ23, Theorem 1], it was already proven that $\binom{n-i}{k-i}/n^{k-i}$ increases in $i$. Hence
$\frac{\binom{n^{\prime}-i}{k-i}}{n^{k-i}}=\frac{\binom{n^{\prime}-i}{k-i}}{(n^{\prime})^{k-i}}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{n}}\right)^{k-i}$
is the product of two expressions both increasing in $i$, and hence is increasing as well. ∎
Now we state and prove our result on the time to see first inner Pareto optimum.
Lemma 6.
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem. Then after at most $e\mu(mk/2)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
Proof.
Denote by $f=(f_{1},\dots,f_{m})$ the $m$OJZJ function. This lemma is already proven if the initial population has at least one inner Pareto optimum. Now we consider that the initial population $P$ contains no inner Pareto optimum. For any $x\in\{0,1\}^{n},$ let $I(x)=\{i\in[1..m/2]\mid|x_{B_{i}}|_{1}\in[k..n^{\prime}-k]\}$ and $L(x)=|I(x)|$. Let $L(P)=\max\{L(x)\mid x\in P\}$. Then $L(P)\geq 0$ and we consider the first time to reach a population $P$ with $L(P)=m/2$, that is, to reach an inner Pareto optimum.
We first show that $L(P)$ cannot decrease. Let $S_{P}=\{x\in P\mid L(x)=L(P)\}$. Let $x\in S_{P}$. From Lemma 4, we know that for the next population, there is $y$ such that $y\succeq x$. Hence, $f_{2i-1}(y)\geq f_{2i-1}(x)$ for any $i\in[1..m/2]$. If $|x_{B_{i}}|_{1}\geq k$, then $f_{2i-1}(x)=k+|x_{B_{i}}|_{1}$. Thus $f_{2i-1}(y)\geq k+|x_{B_{i}}|_{1}\geq 2k$ and hence $|y_{B_{i}}|_{1}\geq k$ from the definition of mOJZJ. An analogous argument based on $f_{2i}$ shows that $|x_{B_{i}}|_{1}\leq n^{\prime}-k$ implies $|y_{B_{i}}|_{1}\leq n^{\prime}-k$. Hence $L(y)\geq L(x)$, which shows that $L(P)$ does not decrease.
Now we consider the increase of $L(P)$ when $L(P)<m/2$. Let $x\in P$ with $L(x)=L(P)$. The probability to select $x$ as a parent is $1/\mu$. For a specific $i\in[1..m/2]\setminus I(x)$, we know that $|x_{B_{i}}|_{1}\in[0..k-1]\cup[n-k+1..n]$. W.l.o.g., let $|x_{B_{i}}|_{1}\in[0..k-1]$. Then the probability of generating an individual $y$ with $|y_{B_{i}}|\in[k..n^{\prime}-k]$ from $x$ is at least
$\binom{n^{\prime}-|x_{B_{i}}|_{1}}{k-|x_{B_{i}}|_{1}}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-|x_{B_{i}}|_{1}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(k-|x_{B_{i}}|_{1})}\geq\binom{n^{\prime}}{k}\frac{1}{n^{k}}\frac{1}{e}\geq\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{k}}\right)^{k}\frac{1}{en^{k}}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{2}{mk}}\right)^{k},$ (1)
where the first inequality is from Lemma 5, and the last equality uses $n^{\prime}=2n/m$. Note that $|[1..m/2]\setminus I(x)|=m/2-L(P)$. Hence, the probability of increasing $L(P)$ in one iteration is at least
$\frac{m/2-L(P)}{\mu}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{2}{mk}}\right)^{k},$
and thus the expected number of iterations to witness such an improvement is at most $e\mu(mk/2)^{k}/(m/2-L(P))$. Hence, the expected number of iterations to generate an inner Pareto optimum is at most
$\sum_{i=0}^{m/2-1}\frac{e\mu}{m/2-i}\mathopen{}\mathclose{{}\left(\frac{mk}{2}}\right)^{k}\leq e\mu\mathopen{}\mathclose{{}\left(\frac{mk}{2}}\right)^{k}(1+\ln m),$
where the last inequality uses $\sum_{j-1}^{m/2}\frac{1}{j}\leq 1+\ln(m/2)\leq 1+\ln m$. ∎
@@
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem. Assume that the current population contains at least one inner Pareto optimum. Then after at most $en\mu(n^{\prime}-2k+1)^{m/2}$ iterations in expectation, all inner Pareto front points are covered.
Proof.
Let $u\in F^{*}$ be an uncovered inner Pareto front point that has some neighbor covered by the population $P$. Since such a neighbor is also a Pareto front point, we know that it is maintained in all future generations. The probability to generate an offspring with function value $u$ is at least $\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n\mu}\geq\frac{1}{en\mu}$. Hence the expected number of iterations to witness such an offspring is at most $en\mu$. From Lemma 4, we know that once such an offspring is generated, $u$ will be covered in all future populations. Noting that there are at most $(n^{\prime}-2k+1)^{m/2}$ inner Pareto front points, we know that after an expected number of at most $en\mu(n^{\prime}-2k+1)^{m/2}$ iterations, all inner Pareto front points are covered. ∎
The last stage is to cover the remaining Pareto front points. The following lemma bounds the runtime of this phase. The key of the proof is that we can divide the Pareto front into several levels and any individual in the $i$-th level can be generated from a point in the $(i-1)$-th level by flipping the right $k$ bits.
@@
Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $e((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})\mu n^{k}$ iterations in expectation, the full Pareto front is covered.
Proof.
We say that a block $x_{B_{i}}$ of an individual $x$ is extreme if $x_{B_{i}}\in\{(0,\dots,0),(1,\dots,1)\}$. We divide the Pareto front points into $m/2+1$ levels $L_{0},L_{1},\dots,L_{m/2}$, where $L_{j}$ consists of the objective values of all Pareto optima having exactly $j$ extreme blocks. Obviously, all inner Pareto front points belong to the 00-th level. Note that any point in $j$-th level ($j\geq 1$) can be generated from an individual having a certain objective value in $L_{j-1}$ via flipping specific $k$ bits, leaving the other bits unchanged. Hence, the probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k}\geq\frac{1}{e\mu n^{k}}.$
From Lemma 4, we know that such a Pareto front point is reached in at most $e\mu n^{k}$ iterations in expectation, and will survive in all future populations. Consider the process of covering all non-inner Pareto front points in the order of increasing level. Since the 00-th level points are already covered from our assumption, and since there are at most $((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})$ non-inner Pareto front points, we know that the remaining levels are covered in at most $e((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})\mu n^{k}$ iterations in expectation. ∎
Summing up the runtimes of the three stages from Lemmas 6 to 8, we obtain the following theorem on the runtime of the full coverage of the Pareto front.
@@
Consider the $m$OJZJ problem with problem size $n$, number of objectives $m$, and gap size $k$. Write $n^{\prime}=2n/m$ for the block length. Assume that $k\leq n^{\prime}/2$. Denote by $M=(n^{\prime}-2k+3)^{m/2}$ the size of the Pareto front and by $\overline{M}$ the size of the largest incomparable set (for which we know that $\overline{M}\leq(2n/m+1)^{m/2}$). Consider using the SMS-EMOA with $\mu\geq\overline{M}$ to solve this problem. Then after at most $e\mu(mk/2)^{k}(1+\ln m)+e\mu Mn^{k}=O(\mu Mn^{k})$ iterations ($\mu+e\mu(mk/2)^{k}(1+\ln m)+e\mu Mn^{k}=O(\mu Mn^{k})$ function evaluations) in expectation, the full Pareto front is covered.
Since $k\leq n^{\prime}/2=n/m$ and $M\geq 2^{m/2}$, we easily see that both runtime expressions in the theorem are $O(\mu Mn^{k})$, even when allowing $k$, $m$ and $\mu$ to depend on $n$.
4.3 Runtime of the GSEMO on $m$OJZJ
Since our arguments above can easily be extended to analyze the runtime of the GSEMO on the $m$OJZJ, we quickly do so for reasons of completeness. The GSEMO is the multiobjective counterpart of the single-objective $(1+1)$ evolutionary algorithm. The initial population of the GSEMO consists of a single randomly generated solution. In each iteration, a solution is picked uniformly at random from the population to generate an offspring via standard bit-wise mutation. If this offspring is not dominated by any solution in the population, it is added to the population and all solutions weakly dominated by it are removed. It is not difficult to see that all solutions in the population of the GSEMO are mutually non-dominated. Hence, the population size of the GSEMO is at most $\overline{M}$ by Lemma 3.